# How Data Mixing Shapes In-Context Learning: Asymptotic Equivalence for Transformers with MLPs

### v1

@@ 4.2 Asymptotic equivalence
Under Assumption 4.6, the entries of $\bm{F}$} are i.i.d. $\mathcal{N}(0,1/\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))))$, ensuring that the components of ${\bm{F}}\text{vec}({\bm{H}}_{{\bm{Z}}})$ have unit variance. Let ${\bm{f}}_{i}$ denote the $i$-th column of $\bm{F}$}. Given the assumptions stated in Section 4.1, we obtain
${\bm{f}}_{i}^{T}\text{vec}({\bm{H}}_{\bm{Z}})\to\mathcal{N}(0,1)\textit{ almost surely,}$ (11)
@@
Set $t:=\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))$. Given ${\bm{H}}_{\bm{Z}}$, we have
@@
The gradient $\bm{G}$} defined in (7) admits the following decomposition: ${\bm{G}}={\bm{u}}{\bm{v}}^{T}+\bm{\Delta}$, where ${\bm{u}}:=\alpha{\bm{w}}$ for some $\alpha\in\mathbb{R}$ and ${\bm{v}}:=\tilde{{\bm{H}}}^{T}\tilde{{\bm{y}}}/(m\sqrt{k})$. Here, ${\bm{u}}{\bm{v}}^{T}$ is the dominant rank-one term while $\|\bm{\Delta}\|$ is the negligible residual term.
@@
Using high-probability tail bounds, this can be shown similarly to the decompositions of the gradient matrix for simpler neural network settings in the literature ba2022highdimensional ; demir2025asymptotic . See Appendix C. ∎
With these components established, we can now leverage the preceding results in conjunction with existing asymptotic analyses of two-layer neural networks demir2025asymptotic to study the in-context learning error defined in (8) for Transformers equipped with nonlinear MLPs. The following theorem identifies a polynomial model that is asymptotically equivalent to such a Transformer architecture.
@@
Theorem 4.12 establishes an asymptotic equivalence between the ICL behavior of a Transformer with an MLP and that of a polynomial model in terms of the ICL errors. This result is significant for several reasons. Firstly, 13 admits a much simpler analysis, allowing precise characterization of ICL error behavior of the Transformer under data mixing and feature learning. Second, it sheds light on the function class learned by the Transformer—specifically, that it effectively learns a low-degree polynomial approximation of the task function. This equivalence opens the door to optimizing nonlinearities in MLPs (inside the Transformer) through polynomial surrogate analysis. These insights extend beyond classical intuition and align with recent trends in Gaussian equivalence literature hu2022universality ; moniri2023theory ; demir2025asymptotic . Note that although similar equivalence results exist for supervised learning by two-layer neural networks moniri2023theory ; demir2025asymptotic , this work represents a cornerstone in adapting those results to a novel setting involving ICL by Transformers with MLPs.
@@
($n=0.5d^{2}$ and $l=d$)
Figure 1: Effects of sample size, context length, and hidden dimension on the ICL errors for linear Transformer (4), Transformer with a nonlinear MLP (5), and the equivalent model (13). The number of data sources is $\mathcal{S}=2$ with equal probability: $\mathbb{P}(s=0)=\mathbb{P}(s=1)=1/2$. For the input vectors, ${\bm{\mu}}_{x,s}=\mathbf{0}$ and ${\bm{\Sigma}}_{x,s}={\bm{I}}_{d}$ for all $s$. For the task vectors, ${\bm{\mu}}_{\xi,s}=\mathbf{0}$ for all $s$ while ${\bm{\Sigma}}_{\xi,0}={\bm{I}}$ and ${\bm{\Sigma}}_{\xi,1}={\bm{I}}_{d}+\theta\bm{\gamma}\bm{\gamma}^{T}$ for some $\theta\asymp d^{2}$ and $\bm{\gamma}\in\mathbb{R}^{d}$ with $\|\bm{\gamma}\|_{2}=1$. The target function $\phi_{s}$ is ReLU, and the target noise is $\Delta_{s}=0.01$ for all $s$. The Transformer is used with two different activation functions: ReLU and tanh. Here, $d=80$, $\eta\asymp d^{2}$, and $\lambda=5\times 10^{-5}$. The degree of the equivalent polynomial model $p$ is set to 44. The average over 20 Monte Carlo runs is plotted.
@@ 4.3.2 Effect of model dimensions
    The nonlinear MLP Transformer (5) outperforms the linear one (4) in terms of ICL error (8). Although outperforming the linear Transformer may sound trivial, without the training of the first layer $\bm{F}$} as in (6), there exist many cases where the Transformer with the nonlinear MLP head fails to outperform the linear one demir2025MLSP . Therefore, outperforming the linear one is a reasonable baseline and an indicator of effective feature learning.
@@
These findings underscore both the expressiveness benefits of nonlinear MLPs and the theoretical significance of our equivalence framework. Additionally, we observe a characteristic double descent phenomenon doubledescent2019 in the ICL error: with respect to sample size in Figure 1(a) and hidden dimension in Figure 1(c). The effect of increasing context length $\ell$ is shown in Figure 1(b), revealing a consistent reduction in ICL error, reflecting improved task estimation with longer contexts.
@@
Figure 2: Data mixing effects on the ICL error: performance of Transformers with ReLU activation (denoted with diamonds (5)), and the equivalent model (denoted with lines (13)) are illustrated for different mixing ratios $\rho$. For this figure, $\rho$ controls the data source mixture for the training, as we define $\mathbb{P}(s=0):=1-\rho$ and $\mathbb{P}(s=1):=\rho$, while the ICL error is calculated as an average over data sources (8). Here, $d=80$, $l=d$, $n=k=0.5d^{2}$, $\lambda=5\times 10^{-5}$, $\eta\asymp d^{2}$ and the target function $\phi_{s}$ is ReLU for all $s$. We initially consider the following setting: for the input vectors, ${\bm{\mu}}_{x,s}=\mathbf{0}$ and ${\bm{\Sigma}}_{x,s}={\bm{I}}_{d}$ for all $s$; for the task vectors, ${\bm{\mu}}_{\xi,s}=\mathbf{0}$ and ${\bm{\Sigma}}_{\xi,s}={\bm{I}}$ for all $s$; and the target noise is $\Delta_{s}=0.01$ for all $s$. For each subfigure, we modify one data property while keeping the rest same and show its effect on the ICL error: in (a), we focus on the input covariance of the second data source by using ${\bm{\Sigma}}_{x,1}={\bm{I}}_{d}+\theta_{x}\bm{\gamma}_{x}\bm{\gamma}_{x}^{T}$ such that $\|{\bm{\Sigma}}_{x,1}\|$ is changed by varying $\theta_{x}$; in (b), we concentrate on the task covariance of the second data source by using ${\bm{\Sigma}}_{\xi,1}={\bm{I}}_{d}+\theta_{\xi}\bm{\gamma}_{\xi}\bm{\gamma}_{\xi}^{T}$ such that $\|{\bm{\Sigma}}_{\xi,1}\|$ is changed by varying $\theta_{\xi}$; and in (c), we modify the noise of the second data source $\Delta_{1}$ while the noise of the first source $\Delta_{0}$ is set to $0.2$ and fixed. We set the degree of the equivalent polynomial model to $p=5$. The average of 20 Monte Carlo runs is plotted.
@@ 4.3.3 Impact of data mixing
We next explore how mixing structured and unstructured data sources affects ICL performance, shown in Figure 2. In these experiments, we vary the second data source by introducing structure into its input covariance (a), task covariance (b), or noise variance (c), while keeping the first source fixed.
@@ 4.3.4 Feature learning with data mixing
    In (a), increasing $\eta$ does not improve ICL error, consistent with our intuition that the first-layer updates do not capture a useful signal when task vectors are isotropic. Here, the intuition comes from the fact that the feature learning enables MLPs to align with task vectors (ba2022highdimensional, , Section 3.2).
@@
Figure 3: Feature learning with data mixing on synthetic and real-world data: the effect of different step sizes $\eta$ is illustrated. For the synthetic data scenarios in (a) and (b), we start with the same initial setting as Figure 2. In (a), we just modify the covariance of the input vectors from the second data source, while we only change the covariance of the task vectors from the second data source in (b). In each case, we add an additional rank-one structure to the covariance, which is the same as Figure 2. We plot the average of 20 Monte Carlo trials for (a) and (b). For the real-world data scenario in (c), we focus on the effect of feature learning on ICL errors for multilingual sentiment analysis using the Multilingual Amazon Reviews Corpus keung2020multilingual . This dataset contains customer reviews (with text and star ratings) in multiple languages, such as English and German. By treating English and German reviews as two distinct data sources, we can vary the mixing ratio across languages, allowing us to evaluate our framework in a more realistic setting. We consider English reviews as source 1 and German reviews as source 0, so a mixing ratio of $\rho=1$ corresponds to entirely English data, and $\rho=0$ to entirely German data. For labels $\{y_{i}\}$, the review star ratings are demeaned and scaled to lie in the range [1,1][-1,1], making the task regression-like. For inputs $\{{\bm{x}}_{i}\}$, each review text is embedded using the multilingual text embedding model called “multilingual-e5-small” wang2024multilingual , the generated 384-dimensional embeddings are reduced to 64 dimensions via PCA (principal component analysis), and then normalized. We group $l$ input-label pairs (of the same language) together to form a context so that we make the problem compatible with our ICL setting. The rest of the details for (c) are $d=l=64$, $n=k=0.25d^{2}$, and $\lambda=5\times 10^{-5}$. The degree of the equivalent polynomial model is set to 55. The mean of 100 Monte Carlo trials is illustrated in (c).
@@ 4.3.5 Data mixing in a real-world scenario: multilingual sentiment analysis
For a real-world example, we consider multilingual sentiment analysis using the Multilingual Amazon Reviews Corpus keung2020multilingual , which contains customer reviews in multiple languages. Treating each language as a distinct source, we examine feature learning under data mixing, as illustrated in Figure 3(c). Consistent with our theoretical predictions, the ICL errors of the Transformer model closely align with those of the equivalent model, even in this multilingual real-world setting. We observe a clear trend (similar to Figure 3(b)) whereby increasing the step size $\eta$ reduces the ICL error, indicating effective feature learning. Finally, a high proportion of English reviews lead to increased performance, consistent with the known strength of the embedding model (used for text) on English data.
@@ 5 Conclusion
We acknowledge that this work is supported partially by TÜBİTAK under project 124E063 in ARDEB 1001 program. We extend our gratitude to TÜBİTAK for its support. S.D. is supported by an AI Fellowship provided by KUIS AI Center and a PhD Scholarship (BİDEB 2211) from TÜBİTAK.
@@ Appendix B Asymptotic distribution of ${\bm{F}}\text{vec}({\bm{H}}_{\bm{Z}})$
Let ${\bm{f}}_{i}$ denote the $i$-th row of the matrix $\bm{F}$}, and define $t:=\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))$. For the purpose of this analysis, we initially assume that ${\bm{H}}_{\bm{Z}}$ is fixed. Under Assumption 4.6, each row ${\bm{f}}_{i}$ is drawn independently from a Gaussian distribution $\mathcal{N}(\mathbf{0},{\bm{I}}/t)$. Consequently, we have the projection
@@
Thus, the task of proving Lemma 4.9 reduces to showing that $\|\text{vec}({\bm{H}}_{{\bm{Z}}})\|_{2}^{2}$ concentrates around $1$, which would imply that ${\bm{f}}_{i}^{T}\text{vec}({\bm{H}}_{{\bm{Z}}})\to\mathcal{N}(0,1)$ almost surely.
@@
$\left|\|{\bm{x}}_{\ell+1}\|^{2}-\text{Tr}(\text{Cov}({\bm{x}}_{\ell+1}))\right|/\text{Tr}(\text{Cov}({\bm{x}}_{\ell+1}))$ $=o(1),$ (S8)
$\left|\|{\bm{b}}\|^{2}-\text{Tr}(\text{Cov}({\bm{b}}))\right|/\text{Tr}(\text{Cov}({\bm{b}}))$ $=o(1).$ (S9)
@@
We can express $\bm{b}$} as: ${\bm{b}}=\frac{1}{\ell}\sum_{i\leq\ell}{\bm{b}}_{i},\quad\text{where}\quad{\bm{b}}_{i}:=\begin{bmatrix}y_{i}{\bm{x}}_{i}\\ y_{i}^{2}\end{bmatrix}$. Since $\phi_{s}$ is Lipschitz and ${\bm{x}}_{i}$ is sub-Gaussian, the composition $y_{i}$ is sub-Gaussian with a constant norm due to Gaussian concentration of Lipschitz functions [40, Theorem 5.2.2].
By Assumption 4.2, we have $\|{\bm{\Sigma}}_{x,s}\|_{2}^{2}=\mathcal{O}(d)$ and $\text{Tr}({\bm{\Sigma}}_{x,s})\asymp d$, so ${\bm{x}}_{i}$ is sub-Gaussian with norm $\mathcal{O}(d^{1/4})$. Then, since the product of sub-Gaussian random variables is sub-exponential [40, Lemma 2.7.7], the term $y_{i}{\bm{x}}_{i}$ is sub-exponential with norm $\mathcal{O}(d^{1/4})$, and $y_{i}^{2}$ is sub-exponential with constant norm. Thus, ${\bm{b}}_{i}$ is a sub-exponential vector with norm $\mathcal{O}(d^{1/4})$.
Applying the (vector version of) Bernstein’s inequality [40, Corollary 2.8.3] to the vector $\bm{b}$} (an average over $l$ samples of a sub-exponential) implies that the deviation $({\bm{b}}-\mathbb{E}[{\bm{b}}])$ has sub-exponential tails with norm $\mathcal{O}(d^{1/4}/\ell)$. Under Assumption 4.1, $\ell/d\in\mathbb{R}$, and hence the norm becomes $\mathcal{O}(d^{-3/4})$.
Finally, applying the Hanson-Wright inequality [40, Theorem 6.2.1] yields the bound in (S8). For (S9), we rely on an extension of the Hanson-Wright inequality to sub-exponential vectors [19]. ∎
@@
$\frac{\left|\|\text{vec}({\bm{H}}_{\bm{Z}})\,|\,(s,\bm{\xi}|s)\|^{2}-\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})\,|\,(s,\bm{\xi}|s)))\right|}{\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})\,|\,(s,\bm{\xi}|s)))}=\left|\frac{\|\text{vec}({\bm{H}}_{\bm{Z}})\,|\,(s,\bm{\xi}|s)\|^{2}}{t}-1\right|=o(1),$
where $t:=\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})))$.
@@
$\text{Cov}(\text{vec}({\bm{H}}_{\bm{Z}}))$ $=\mathbb{E}[({\bm{b}}\otimes{\bm{x}}_{\ell+1})({\bm{b}}\otimes{\bm{x}}_{\ell+1})^{T}]$ (S10)
$=\mathbb{E}[({\bm{b}}\otimes{\bm{x}}_{\ell+1})({\bm{b}}^{T}\otimes{\bm{x}}_{\ell+1}^{T})]$ (S11)
$=\mathbb{E}[({\bm{b}}{\bm{b}}^{T})\otimes({\bm{x}}_{\ell+1}{\bm{x}}_{\ell+1}^{T})]$ (S12)
$=\text{Cov}({\bm{b}})\otimes\text{Cov}({\bm{x}}_{\ell+1}),$ (S13)
@@
$\|\text{vec}({\bm{H}}_{\bm{Z}})\|^{2}=\|{\bm{b}}\otimes{\bm{x}}_{\ell+1}\|^{2}=\|{\bm{b}}\|^{2}\cdot\|{\bm{x}}_{\ell+1}\|^{2},$ (S14)
@@
$\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})))=\text{Tr}(\text{Cov}({\bm{b}}))\cdot\text{Tr}(\text{Cov}({\bm{x}}_{\ell+1})).$ (S15)

### v2

@@ 4.2 Asymptotic equivalence
Under Assumption 4.6, the entries of $\bm{F}$} are i.i.d. $\mathcal{N}(0,1/\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))))$, ensuring that the components of ${\bm{F}}\text{vec}({\bm{H}}_{{\bm{Z}}})$ have unit variance. Let ${\bm{f}}_{i}$ denote the $i$-th row of $\bm{F}$}. Given the assumptions stated in Section 4.1, we obtain
${\bm{f}}_{i}^{T}\text{vec}({\bm{H}}_{\bm{Z}})\to\mathcal{N}(0,1)\textit{ converges in distribution,}$ (11)
@@
Set $t:=\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))$. Given ${\bm{H}}_{\bm{Z}}$, we have
@@
The gradient $\bm{G}$} defined in (7) admits the following decomposition: ${\bm{G}}={\bm{u}}{\bm{v}}^{T}+\bm{\Delta}$, where ${\bm{u}}:=\alpha{\bm{w}}$ for some $\alpha\in\mathbb{R}$ and ${\bm{v}}:=\tilde{{\bm{H}}}^{T}\tilde{{\bm{y}}}/(m\sqrt{k})$. Here, ${\bm{u}}{\bm{v}}^{T}$ is the dominant rank-one term while $\|\bm{\Delta}\|_{2}$ is the negligible residual term.
@@
Using high-probability tail bounds, this can be shown similarly to the decompositions of the gradient matrix for simpler neural network settings in the literature Ba et al. (2022); Demir and Dogan (2025a). See Appendix C. ∎
With these components established, we can now leverage the preceding results in conjunction with existing asymptotic analyses of two-layer neural networks Demir and Dogan (2025a) to study the in-context learning error defined in (8) for Transformers equipped with nonlinear MLPs. The following theorem identifies a polynomial model that is asymptotically equivalent to such a Transformer architecture.
@@
Theorem 4.12 establishes an asymptotic equivalence between the ICL behavior of a Transformer with an MLP and that of a polynomial model in terms of the ICL errors. This result is significant for several reasons. Firstly, (13) admits a much simpler analysis, allowing precise characterization of ICL error behavior of the Transformer under data mixing and feature learning. Second, it sheds light on the function class learned by the Transformer—specifically, that it effectively learns a low-degree polynomial approximation of the task function. This equivalence opens the door to optimizing nonlinearities in MLPs (inside the Transformer) through polynomial surrogate analysis. These insights extend beyond classical intuition and align with recent trends in Gaussian equivalence literature Hu and Lu (2023); Moniri et al. (2024); Demir and Dogan (2025a). Note that although similar equivalence results exist for supervised learning by two-layer neural networks Moniri et al. (2024); Demir and Dogan (2025a), this work represents a cornerstone in adapting those results to a novel setting involving ICL by Transformers with MLPs.
@@
($n=0.5d^{2}$ and $\ell=d$)
Figure 1: Effects of sample size, context length, and hidden dimension on the ICL errors for linear Transformer (4), Transformer with a nonlinear MLP (5), and the equivalent model (13). The number of data sources is $\mathcal{S}=2$ with equal probability: $\mathbb{P}(s=0)=\mathbb{P}(s=1)=1/2$. For the input vectors, ${\bm{\mu}}_{x,s}=\mathbf{0}$ and ${\bm{\Sigma}}_{x,s}={\bm{I}}_{d}$ for all $s$. For the task vectors, ${\bm{\mu}}_{\xi,s}=\mathbf{0}$ for all $s$ while ${\bm{\Sigma}}_{\xi,0}={\bm{I}}_{d}$ and ${\bm{\Sigma}}_{\xi,1}={\bm{I}}_{d}+\theta\bm{\gamma}\bm{\gamma}^{T}$ for some $\theta\asymp d^{2}$ and $\bm{\gamma}\in\mathbb{R}^{d}$ with $\|\bm{\gamma}\|_{2}=1$. The target function $\phi_{s}$ is ReLU, and the target noise is $\Delta_{s}=0.01$ for all $s$. The Transformer is used with two different activation functions: ReLU and tanh. Here, $d=80$, $\eta\asymp d^{2}$, and $\lambda=5\times 10^{-5}$. The degree of the equivalent polynomial model $p$ is set to 44. The average over 20 Monte Carlo runs is plotted.
@@ 4.3.2 Effect of model dimensions
    The nonlinear MLP Transformer (5) outperforms the linear one (4) in terms of ICL error (8). Although outperforming the linear Transformer may sound trivial, without the training of the first layer $\bm{F}$} as in (6), there exist many cases where the Transformer with the nonlinear MLP head fails to outperform the linear one Demir and Dogan (2025b). Therefore, outperforming the linear one is a reasonable baseline and an indicator of effective feature learning.
@@
These findings underscore both the benefits of nonlinear MLPs and the theoretical significance of our equivalence framework. Additionally, we observe a characteristic double descent phenomenon Belkin et al. (2019) in the ICL error: with respect to sample size in Figure 1(a) and hidden dimension in Figure 1(c). The effect of increasing context length $\ell$ is shown in Figure 1(b), revealing a consistent reduction in ICL error, reflecting improved task estimation with longer contexts.
@@
Figure 2: Data mixing effects on the ICL error: performance of Transformers with ReLU activation (denoted with diamonds (5)), and the equivalent model (denoted with lines (13)) are illustrated for different mixing ratios $\rho$. For this figure, $\rho$ controls the data source mixture for the training, as we define $\mathbb{P}(s=0):=1-\rho$ and $\mathbb{P}(s=1):=\rho$, while the ICL error is calculated as an average over data sources (8). Here, $d=80$, $l=d$, $n=k=0.5d^{2}$, $\lambda=5\times 10^{-5}$, $\eta\asymp d^{2}$ and the target function $\phi_{s}$ is ReLU for all $s$. We initially consider the following setting: for the input vectors, ${\bm{\mu}}_{x,s}=\mathbf{0}$ and ${\bm{\Sigma}}_{x,s}={\bm{I}}_{d}$ for all $s$; for the task vectors, ${\bm{\mu}}_{\xi,s}=\mathbf{0}$ and ${\bm{\Sigma}}_{\xi,s}={\bm{I}}_{d}$ for all $s$; and the target noise is $\Delta_{s}=0.01$ for all $s$. For each subfigure, we modify one data property while keeping the rest same and show its effect on the ICL error: in (a), we focus on the input covariance of the second data source by using ${\bm{\Sigma}}_{x,1}={\bm{I}}_{d}+\theta_{x}\bm{\gamma}_{x}\bm{\gamma}_{x}^{T}$ such that $\|{\bm{\Sigma}}_{x,1}\|_{2}$ is changed by varying $\theta_{x}$; in (b), we concentrate on the task covariance of the second data source by using ${\bm{\Sigma}}_{\xi,1}={\bm{I}}_{d}+\theta_{\xi}\bm{\gamma}_{\xi}\bm{\gamma}_{\xi}^{T}$ such that $\|{\bm{\Sigma}}_{\xi,1}\|_{2}$ is changed by varying $\theta_{\xi}$; and in (c), we modify the noise of the second data source $\Delta_{1}$ while the noise of the first source $\Delta_{0}$ is set to $0.2$ and fixed. We set the degree of the equivalent polynomial model to $p=5$. The average of 20 Monte Carlo runs is plotted.
@@ 4.3.3 Impact of data mixing
We next explore how mixing structured and unstructured data sources affects ICL performance, shown in Figure 2. In these experiments, we vary the second data source by modifying its input covariance (a), task covariance (b), or noise variance (c), while keeping the first source fixed.
@@ 4.3.4 Feature learning with data mixing
    In (a), increasing $\eta$ does not improve ICL error, consistent with our intuition that the first-layer updates do not capture a useful signal when task vectors are isotropic. Here, the intuition comes from the fact that the feature learning enables MLPs to align with task vectors (Ba et al., 2022, Section 3.2).
@@
Figure 3: Feature learning with data mixing on synthetic and real-world data: the effect of different step sizes $\eta$ is illustrated. For the synthetic data scenarios in (a) and (b), we start with the same initial setting as Figure 2. In (a), we just modify the covariance of the input vectors from the second data source, while we only change the covariance of the task vectors from the second data source in (b). In each case, we add an additional rank-one structure to the covariance, which is the same as Figure 2. We plot the average of 20 Monte Carlo trials for (a) and (b). For the real-world data scenario in (c), we focus on the effect of feature learning on ICL errors for multilingual sentiment analysis using the Multilingual Amazon Reviews Corpus Keung et al. (2020). This dataset contains customer reviews (with text and star ratings) in multiple languages, such as English and German. By treating English and German reviews as two distinct data sources, we can vary the mixing ratio across languages, allowing us to evaluate our framework in a more realistic setting. We consider English reviews as source 1 and German reviews as source 0, so a mixing ratio of $\rho=1$ corresponds to entirely English data, and $\rho=0$ to entirely German data. For labels $\{y_{i}\}$, the review star ratings are demeaned and scaled to lie in the range [1,1][-1,1], making the task regression-like. For inputs $\{{\bm{x}}_{i}\}$, each review text is embedded using the multilingual text embedding model called “multilingual-e5-small” Wang et al. (2024), the generated 384-dimensional embeddings are reduced to 64 dimensions via PCA (principal component analysis), and then normalized. We group $l$ input-label pairs (of the same language) together to form a context so that we make the problem compatible with our ICL setting. The rest of the details for (c) are $d=l=64$, $n=k=0.25d^{2}$, and $\lambda=5\times 10^{-5}$. The degree of the equivalent polynomial model is set to 55. The mean of 100 Monte Carlo trials is illustrated in (c).
@@ 4.3.5 Data mixing in a real-world scenario: multilingual sentiment analysis
For a real-world example, we consider multilingual sentiment analysis using the Multilingual Amazon Reviews Corpus Keung et al. (2020), which contains customer reviews in multiple languages. Treating each language as a distinct source, we examine feature learning under data mixing, as illustrated in Figure 3(c). Consistent with our theoretical predictions, the ICL errors of the Transformer model closely align with those of the equivalent model, even in this multilingual real-world setting. We observe a clear trend (similar to Figure 3(b)) whereby increasing the step size $\eta$ reduces the ICL error, indicating effective feature learning. Finally, a higher proportion of English reviews can lead to increased performance, consistent with the known strength of the embedding model (for text) on English data.
@@ 5 Conclusion
We acknowledge that this work is supported partially by TÜBİTAK under project 124E063 in ARDEB 1001 program. We extend our gratitude to TÜBİTAK for its support. S.D. is supported by an AI Fellowship provided by KUIS AI Center and a PhD Scholarship (BİDEB 2211) from TÜBİTAK. Z.D. is also partially supported by the BİDEB 2224-A program from TÜBİTAK.
@@ Appendix B Asymptotic distribution of ${\bm{F}}\text{vec}({\bm{H}}_{\bm{Z}})$
Let ${\bm{f}}_{i}$ denote the $i$-th row of the matrix $\bm{F}$}, and define $t:=\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))$. For the purpose of this analysis, we initially assume that ${\bm{H}}_{\bm{Z}}$ is fixed. Under Assumption 4.6, each row ${\bm{f}}_{i}$ is drawn independently from a Gaussian distribution $\mathcal{N}(\mathbf{0},{\bm{I}}_{d(d+1)}/t)$. Consequently, we have
@@
Thus, the task of proving Lemma 4.9 reduces to showing that $\|\text{vec}({\bm{H}}_{{\bm{Z}}})\|_{2}^{2}/t$ concentrates around $1$, which would imply that ${\bm{f}}_{i}^{T}\text{vec}({\bm{H}}_{{\bm{Z}}})\to\mathcal{N}(0,1)$.
@@
$\left|\|{\bm{x}}_{\ell+1}\|_{2}^{2}-\operatorname{Tr}(\mathrm{Cov}({\bm{x}}_{\ell+1}))\right|/\operatorname{Tr}(\mathrm{Cov}({\bm{x}}_{\ell+1}))$ $=o(1),$ (S8)
$\left|\|{\bm{b}}\|_{2}^{2}-\operatorname{Tr}(\mathbb{E}[{\bm{b}}{\bm{b}}^{T}])\right|/\operatorname{Tr}(\mathbb{E}[{\bm{b}}{\bm{b}}^{T}])$ $=o(1).$ (S9)
@@
We can express $\bm{b}$} as: ${\bm{b}}=\frac{1}{\ell}\sum_{i\leq\ell}{\bm{b}}_{i},\quad\text{where}\quad{\bm{b}}_{i}:=\begin{bmatrix}y_{i}{\bm{x}}_{i}\\ y_{i}^{2}\end{bmatrix}$. Since $\phi_{s}$ is Lipschitz and ${\bm{x}}_{i}$ is sub-Gaussian, the composition $y_{i}$ is sub-Gaussian with a constant norm due to Gaussian concentration of Lipschitz functions [39, Theorem 5.2.2].
By Assumption 4.2, we have $\|{\bm{\Sigma}}_{x,s}\|_{2}^{2}=\mathcal{O}(d)$ and $\operatorname{Tr}({\bm{\Sigma}}_{x,s})\asymp d$, so ${\bm{x}}_{i}$ is sub-Gaussian with norm $\mathcal{O}(d^{1/4})$. Then, since the product of sub-Gaussian random variables is sub-exponential [39, Lemma 2.7.7], the term $y_{i}{\bm{x}}_{i}$ is sub-exponential with norm $\mathcal{O}(d^{1/4})$, and $y_{i}^{2}$ is sub-exponential with constant norm. Thus, ${\bm{b}}_{i}$ is a sub-exponential vector with norm $\mathcal{O}(d^{1/4})$.
Applying the (vector version of) Bernstein’s inequality [39, Corollary 2.8.3] to the vector $\bm{b}$} (an average over $l$ samples of a sub-exponential) implies that the deviation $({\bm{b}}-\mathbb{E}[{\bm{b}}])$ has sub-exponential tails with norm $\mathcal{O}(d^{1/4}/\ell)$. Under Assumption 4.1, $\ell/d\in\mathbb{R}$, and hence the norm becomes $\mathcal{O}(d^{-3/4})$.
Finally, applying the Hanson-Wright inequality [39, Theorem 6.2.1] yields the bound in (S8). For (S9), we rely on an extension of the Hanson-Wright inequality to sub-exponential vectors [19]. ∎
@@
$\frac{\left|\|\text{vec}({\bm{H}}_{\bm{Z}})\,|\,(s,\bm{\xi}|s)\|_{2}^{2}-\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})\,|\,(s,\bm{\xi}|s)))\right|}{\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})\,|\,(s,\bm{\xi}|s)))}=\left|\frac{\|\text{vec}({\bm{H}}_{\bm{Z}})\,|\,(s,\bm{\xi}|s)\|_{2}^{2}}{t}-1\right|=o(1),$
where $t:=\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})))$.
@@
$\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}}))$ $=\mathbb{E}[({\bm{b}}\otimes{\bm{x}}_{\ell+1})({\bm{b}}\otimes{\bm{x}}_{\ell+1})^{T}],$ (S10)
$=\mathbb{E}[({\bm{b}}\otimes{\bm{x}}_{\ell+1})({\bm{b}}^{T}\otimes{\bm{x}}_{\ell+1}^{T})],$ (S11)
$=\mathbb{E}[({\bm{b}}{\bm{b}}^{T})\otimes({\bm{x}}_{\ell+1}{\bm{x}}_{\ell+1}^{T})],$ (S12)
$=\mathbb{E}[{\bm{b}}{\bm{b}}^{T}]\otimes\mathrm{Cov}({\bm{x}}_{\ell+1}),$ (S13)
@@
$\|\text{vec}({\bm{H}}_{\bm{Z}})\|_{2}^{2}=\|{\bm{b}}\otimes{\bm{x}}_{\ell+1}\|_{2}^{2}=\|{\bm{b}}\|_{2}^{2}\cdot\|{\bm{x}}_{\ell+1}\|_{2}^{2},$ (S14)
@@
$\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})))=\operatorname{Tr}(\mathbb{E}[{\bm{b}}{\bm{b}}^{T}])\cdot\operatorname{Tr}(\mathrm{Cov}({\bm{x}}_{\ell+1})).$ (S15)